A 10.0 cm length of copper wire has a uniform cross-sectional area of 1.0 mm2. If the drift speed of electrons in the wire is: 2.0x10-4 m/s and the electrons all move directly along the wire, what is the magnitude of the current density? What is the total current carried by the wire.
Answer #1

a)

drift speed,

Vd = J/(n*q)

J --> current density = ?

n --> charge carrier density = 8.5*10^28 /m^3

q --> charge of electron = 1.6*10^-19 C

so,

J = vd*n*q

= 2*10^-4*8.5*10^28*1.6*10^-19

= 2.72*10^6 A/m^2

b) current through wire, I = J*A

= 2.72*10^6*(1*10^-3)^2

= 2.72 A